A 10-kg duffle bag is undergoing a vertical acceleration of 0.65 m/s2 (positive for up, negative for down), while its owner is holding it with an upward force. What is the magnitude of this force?
1
Expert's answer
2011-02-24T06:06:14-0500
For upward acceleration: a-g = 0.65 ;& g=9.8 , thus& a = 9.8 + 0.65 = 10.45. The force would be F = ma = 10*10.45 = 104.5 N.
For downward: a-g = -0.65, a = 9.8 - 0.65 = 9.15, The force would be F = 10* 9.15 = 91.5 N.